Time Frequency Brown Bag Seminar

Wednesday, April 8, 1998

12:00 noon

EQuad E415

Speaker: Dr. Hamid Jafarkhani, AT&T Labs - Research

Title: New trellis coded quantization schemes

Abstract:

A powerful source coding scheme for memoryless sources is trellis coded quantization. The performance of a trellis coded quantizer (TCQ) is much better than that of the best scalar quantizer (Lloyd-Max quantizer) at the same rate. This is due to the fact that granular gain (space-filling advantage) can be achieved by using a TCQ. TCQ has been used in many speech and image coding systems. Using a tensor product trellis, we show how to design successively refinable TCQs and multiple description TCQs.

In recent years, successively refinable or rate-scalable source coders have received growing attention. An embedded output bit stream is provided by successively refinable source coders, i.e., by selecting different sub-streams of the output, various levels of encoding rate and

distortion can be achieved. Successive refinability enables progressive transmission which is an essential component of image and multimedia browsing systems. Unfortunately the original TCQ proposed by Marcellin and Fischer does not lend itself to successive refinability. In this work, we present a new structure for TCQs which is inherently rate-scalable and is called successively refinable TCQ.

In a multiple description source coding scheme, the source encoder outputs two bit streams which are transmitted over two separate channels. There are three receivers, two of them receive only one of the descriptions (side decoders) and the third one receives both descriptions (central decoder). The objective is to minimize the central distortion, subject to constraints on the side distortions. In this work, we show how to design a multiple description TCQ.